# On Erlang mixture approximations for differential equations

### v2

@@ Abstract
In this paper, we propose a general approach for approximate simulation and analysis of delay differential equations (DDEs) with distributed time delays based on methods for ordinary differential equations (ODEs). The key innovation is that we 1) approximate the kernel by the probability density function of an Erlang mixture and 2) use the linear chain trick to transform the approximate DDEs to ODEs. Furthermore, we prove that an approximation with infinitely many terms converges for continuous and bounded kernels and for specific choices of the coefficients. We show that the approximate ODEs can be used to assess the stability of the steady states of the original DDEs and that the solution to the ODEs converges if the kernel is also exponentially bounded. Additionally, we propose an approach based on bisection and least-squares estimation for determining optimal parameter values in the approximation. Finally, we present numerical examples that demonstrate the accuracy and convergence rate obtained with the optimal parameters and the efficacy of the proposed approach for bifurcation analysis and Monte Carlo simulation. The numerical examples involve a modified logistic equation, chemotherapy-induced myelosuppression, and a point reactor kinetics model of a molten salt nuclear fission reactor.
@@ 1 Introduction
Consequently, it is desirable to identify approaches that can directly take advantage of existing theory, methods, and software for ODEs, and for a specific class of kernels, it is possible to transform DDEs with distributed delays to ODEs. Specifically, the transformation is called the linear chain trick (LCT) [36, 41, 50], and it is applicable when the kernel is given by the probability density function of an Erlang distribution (referred to as an Erlang kernel). See also [12] on sufficient and necessary conditions under which DDEs can be transformed to ODEs. The LCT has been widely used, and Hurtado and Kirosingh [26] and Hurtado and Richards [27, 28] generalized it to transform a wider range of stochastic mean field models to ODEs. Additionally, both Cassidy et al. [9] and Krzyzanski [33] have proposed to simplify the simulation and analysis of specific DDEs by approximating the involved gamma kernels by a hypoexponential kernel and a truncated binomial series, respectively, and Guglielmi and Hairer [18] consider exponential mixture approximations and demonstrate their approach for DDEs involving both gamma and Pareto kernels. However, to the best of the author’s knowledge, there are no general-purpose LCTs or approximations for broader classes of kernels.
@@
The remainder of the paper is structured as follows. In Section 2 and 3, we present the notation used and the system considered in this paper, respectively. Next, in Section 4, we present the mixed Erlang approximation and prove that it converges, and in Section 5, we describe how to use the LCT to transform the approximate system of DDEs to a set of ODEs. In Section 6, we describe the least-squares approach, and in Section 7 and 8, we present numerical examples. Finally, conclusions are presented in Section 9.
@@ 2 Preliminaries
The spaces $\mathbb{R}_{\geq 0}=\{x\in\mathbb{R}|x\geq 0\}$ and $\mathbb{N}_{\geq 0}=\{0,1,\ldots\}$ contain the non-negative real numbers and integers, respectively, and $\mathbb{R}_{>0}=\{x\in\mathbb{R}|x>0\}$ and $\mathbb{R}_{\leq 0}=\{x\in\mathbb{R}|x\leq 0\}$ are the spaces of positive and non-negative real numbers. We denote by $\odot$ the Hadamard (or elementwise) product (also called the Schur product) [22, 23], i.e., if $c=a\odot b$, then $c_{i}=a_{i}b_{i}$. Furthermore, $\det A$ is the determinant of the matrix argument, $A$, and $\operatorname{diag}a$ denotes a diagonal matrix whose diagonal elements are the elements of the vector (or set) argument, $a$, i.e., if $C=\operatorname{diag}a$, then $C_{ii}=a_{i}$ and $C_{ij}=0$ for $i\neq j$. We will make repeated use of the identity
@@ 3 System
where $r:\mathbb{R}\rightarrow\mathbb{R}^{n_{z}}$ is the delayed variable, and each element of $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}^{n_{z}}$ is an exponentially bounded regular kernel (see Definition 3.13.2). Furthermore, $h:\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}^{n_{z}}$ is the memory function. We assume that $f$ and $h$ are differentiable in their arguments and that $r$ is bounded, i.e., that $r(t)\leq K_{r}$ for all $t\in(-\infty,t_{f}]$. We refer to the paper by Ponosov et al. [41, Thm. 1] for more details on the existence and uniqueness of solutions to the initial value problem (2)–(3). See also the book by Hale and Lunel [21].
@@
A scalar-valued kernel, $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, is regular if it satisfies the following properties.
  1. 1.
    It is non-negative and bounded, i.e., $0\leq\alpha(t)\leq K$ for all $t\in\mathbb{R}_{\geq 0}$ and for some finite $K\in\mathbb{R}_{>0}$.
  2. 2.
@@
  3. 3.
    It is normalized such that
(4) $\int\limits_{0}^{\infty}\alpha(t)\,\mathrm{d}t$ $=1.$
@@
A scalar-valued kernel, $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, is exponentially bounded if there exist $L,\rho\in\mathbb{R}_{>0}$ such that
(5) $\alpha(t)$ $\leq Le^{-\rho t}$
@@
For a given system of DDEs with distributed time delays, each element of $\alpha$ may not satisfy (4). However, as they are assumed to be nonzero and non-negative, it is straightforward to normalize them. Next, we present a few well-known corollaries about the steady states of (2b)–(3) and their stability.
@@
(6) $0$ $=f(\bar{x},\bar{z}),$ $\bar{z}$ $=\bar{r}=h(\bar{x}).$
Proof 3.4.
@@
(7) $z(t)$ $=\int_{-\infty}^{t}\alpha(t-s)\odot\bar{r}\,\mathrm{d}s=\int_{-\infty}^{t}\alpha(t-s)\,\mathrm{d}s\odot\bar{r}=\bar{r},$
for all $t$, where we have used the property (4) of each element of $\alpha$.
@@
where $P:\mathbb{C}\rightarrow\mathbb{C}$ is the characteristic function and $I\in\mathbb{R}^{n_{x}\times n_{x}}$ is an identity matrix. The matrices $F\in\mathbb{R}^{n_{x}\times n_{x}}$, $G\in\mathbb{R}^{n_{x}\times n_{z}}$, and $H\in\mathbb{R}^{n_{z}\times n_{x}}$ are the Jacobians of the right-hand side function and the delay function evaluated in the steady state:
Proof 3.6.
The linearized system corresponding to (2b)–(3) describes the evolution of the deviation variable $X:\mathbb{R}\rightarrow\mathbb{R}^{n_{x}}$:
See, e.g., [10, 11, 37] for proofs of the condition (8) for asymptotic stability of the linearized system in (10).
4 Erlang mixture approximation
In this section, we introduce the Erlang kernel and the Erlang mixture kernel, and we prove that the parameters in the latter can be chosen such that it converges to any regular kernel. Specifically, in Section 4.1, we express an infinite-order Erlang mixture kernel in terms of a so-called delta family and derive expressions for the coefficients based on the kernel, $\alpha$, to be approximated. Next, in Section 4.2, we use the delta family to prove pointwise convergence of the Erlang mixture approximation. In Section 4.3, we present a previous convergence result by Tijms [51] which is used in Section 4.4 to prove several properties of Erlang mixture kernels. These properties are used to prove that the integral of the absolute error converges to zero in Section 4.5 and to relate the stability analyses of the original DDEs in Section 3 and the approximate ODEs described in Section 5. Figure 1 shows examples of Erlang kernels and Erlang mixture kernels for different rate parameters and coefficients. For simplicity of the presentation, we only consider scalar kernels in this section.
@@ 4 Erlang mixture approximation
Definition 4.2.
The $M$’th order Erlang mixture kernel, $\hat{\alpha}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, with rate parameter $a\in\mathbb{R}_{>0}$ and coefficients $c_{m}\in[0,1]$ for $m=0,\ldots,M$ is given by
(13) $\hat{\alpha}(t)$ $=\sum_{m=0}^{M}c_{m}\ell_{m}(t),$
where $M\in\mathbb{N}_{\geq 0}$ and $\ell_{m}$ is an $m$’th order Erlang kernel with rate parameter $a$. The coefficients must sum to one:
(14) $\sum_{m=0}^{M}c_{m}$ $=1.$
A straightforward consequence of (12) and (14) is that
(15) $\int_{0}^{\infty}\hat{\alpha}(t)\,\mathrm{d}t$ $=1.$

### v3

@@ Abstract
In this paper, we propose a general approach for approximate simulation and analysis of delay differential equations (DDEs) with distributed time delays based on methods for ordinary differential equations (ODEs). The key innovation is that we 1) propose an Erlang mixture approximation of the kernel in the DDEs and 2) use the linear chain trick to transform the resulting approximate DDEs to ODEs. Furthermore, we prove that the approximation converges for continuous and bounded kernels and for specific choices of the coefficients if the number of terms increases sufficiently fast. We show that the approximate ODEs can be used to assess the stability of the steady states of the original DDEs and that the solution to the ODEs converges if the kernel is also exponentially bounded. Additionally, we propose an approach based on bisection and least-squares estimation for determining optimal parameter values in the approximation. Finally, we present numerical examples that demonstrate the accuracy and convergence rate obtained with the optimal parameters and the efficacy of the proposed approach for bifurcation analysis and Monte Carlo simulation. The numerical examples involve a modified logistic equation, chemotherapy-induced myelosuppression, and a point reactor kinetics model of a molten salt nuclear fission reactor.
@@ 1 Introduction
Consequently, it is desirable to identify approaches that can directly take advantage of existing theory, methods, and software for ODEs, and for a specific class of kernels, it is possible to transform DDEs with distributed delays to ODEs. Specifically, the transformation is called the linear chain trick (LCT) [35, 40, 50], and it is applicable when the kernel is given by the probability density function of an Erlang distribution (referred to as an Erlang kernel). See also [12] on sufficient and necessary conditions under which DDEs can be transformed to ODEs. The LCT has been widely used, and Hurtado and Kirosingh [25] and Hurtado and Richards [26, 27] generalized it to transform a wider range of stochastic mean field models to ODEs. Additionally, both Cassidy et al. [9] and Krzyzanski [32] have proposed to simplify the simulation and analysis of specific DDEs by approximating the involved gamma kernels by a hypoexponential kernel and a truncated binomial series, respectively, and Guglielmi and Hairer [18] consider exponential mixture approximations and demonstrate their approach for DDEs involving both gamma and Pareto kernels. However, to the best of the author’s knowledge, Erlang mixture approximations of general continuous and exponentially bounded kernels have not previously been combined with the LCT to approximate DDEs with distributed time delays by ODEs.
@@
The remainder of the paper is structured as follows. In Section 2 and 3, we present the notation used and the system considered in this paper, respectively. Next, we present the main theoretical results of the paper in Section 4, and in Section 5, we describe the least-squares approach. In Section 6 and 7, we present numerical examples, and conclusions are presented in Section 8.
@@ 2 Preliminaries
The spaces $\mathbb{R}_{\geq 0}=\{x\in\mathbb{R}|x\geq 0\}$ and $\mathbb{N}_{\geq 0}=\{0,1,\ldots\}$ contain the non-negative real numbers and integers, respectively, and $\mathbb{R}_{>0}=\{x\in\mathbb{R}|x>0\}$ and $\mathbb{R}_{\leq 0}=\{x\in\mathbb{R}|x\leq 0\}$ are the spaces of positive and non-positive real numbers. We denote by $\odot$ the Hadamard (or elementwise) product (also called the Schur product) [21, 22], i.e., if $c=a\odot b$, then $c_{i}=a_{i}b_{i}$. Furthermore, $\det A$ is the determinant of the matrix argument, $A$, and $\operatorname{diag}a$ denotes a diagonal matrix whose diagonal elements are the elements of the vector (or set) argument, $a$, i.e., if $C=\operatorname{diag}a$, then $C_{ii}=a_{i}$ and $C_{ij}=0$ for $i\neq j$. We will use the $\infty$-norm defined by
(1) $\|a\|_{\infty}$ $=\max\,\{|a_{i}|\}_{i=1}^{n},$ $\|A\|_{\infty}$ $=\max\,\left\{\sum_{j=1}^{m}|A_{ij}|\right\}_{i=1}^{n}$
for a vector $a\in\mathbb{R}^{n}$ and a matrix $A\in\mathbb{R}^{n\times m}$. If $A=\operatorname{diag}a$ is diagonal, then $\|A\|_{\infty}=\|a\|_{\infty}$. We will make repeated use of the identity
@@
The main theoretical results of the paper are related to convergence [48, Chap. 7]. A function $f_{n}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}$ converges pointwise to $f:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}$ (i.e., $f_{n}(x)\rightarrow f(x)$) as $n\rightarrow\infty$ if there exists an $N:\mathbb{R}_{\geq 0}\rightarrow\mathbb{N}$ for every $\epsilon\in\mathbb{R}_{>0}$ and $x\in\mathbb{R}_{\geq 0}$ such that $|f_{n}(x)-f(x)|<\epsilon$ for all $n>N(x)$. If $N$ can be chosen independently of $x$, the convergence is uniform. We will also show results on weak convergence, i.e., that the integral converges: $\int_{0}^{x}f_{n}(y)\,\mathrm{d}y\rightarrow\int_{0}^{x}f(y)\,\mathrm{d}y$ as $n\rightarrow\infty$. Finally, we will also show convergence of the absolute difference (convergence in $L^{1}$-norm): $\int_{0}^{\infty}|f_{n}(x)-f(x)|\,\mathrm{d}x\rightarrow 0$ as $n\rightarrow\infty$. We will also use the result that if $f_{n}(x)\rightarrow g_{n}(x)$ and $g_{n}(x)\rightarrow f(x)$ as $n\rightarrow\infty$ where $g_{n}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}$, then $f_{n}(x)\rightarrow f(x)$ as well because the error can be bounded through a triangle inequality: $|f_{n}(x)-f(x)|=|f_{n}(x)-g_{n}(x)+g_{n}(x)-f(x)|\leq|f_{n}(x)-g_{n}(x)|+|g_{n}(x)-f(x)|$. Both terms can be made arbitrarily small. In particular, we will use this result together with Stirling’s approximation [46]: $\ln n!\approx n\ln n-n+\frac{1}{2}\ln 2\pi n$. The error of this approximation decreases with $1/n$, i.e., the approximation converges as $n\rightarrow\infty$. The concepts are similar if $n$ is a real number instead of a natural number and if the domain and range of $f_{n}$, $g_{n}$, and $f$ are different.
@@ 3 System
where $r:\mathbb{R}\rightarrow\mathbb{R}^{n_{z}}$ is the delayed variable, and each element of $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{n_{z}}$ is an exponentially bounded regular kernel (see Definition 3.23.3). Furthermore, $h:\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}^{n_{z}}$ is the memory function. We assume that $f$ and $h$ are differentiable in their arguments and that $r$ is bounded, i.e., that $\|r(t)\|_{\infty}\leq K_{r}$ for all $t\in(-\infty,t_{f}]$. Additionally, we assume that $x_{0}$ is continuous and exponentially bounded, i.e., $\|x_{0}(t)\|_{\infty}\leq L_{0}e^{-\sigma t}$, and that the rate parameter, $\sigma\in\mathbb{R}_{>0}$, is strictly smaller than all of the rate parameters in the exponential bounds for the kernels. We refer to the paper by Ponosov et al. [40, Thm. 1] for more details on the existence and uniqueness of solutions to the initial value problem (IVP) (3)–(4). See also the book by Hale and Lunel [20].
@@
For a scalar-valued kernel, $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}$, the integrals $\beta:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}$ and $\rho:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ are defined by
(5) $\beta(t)$ $=\int_{0}^{t}\alpha(s)\,\mathrm{d}s,$ $\rho(t)$ $=\int_{0}^{t}|\alpha(s)|\,\mathrm{d}s.$
Definition 3.2.
A scalar-valued kernel, $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}$, is regular if it satisfies the following conditions.
  1. (a)
    It is bounded, i.e., $|\alpha(t)|\leq K$ for all $t\in\mathbb{R}_{\geq 0}$ and for some finite $K\in\mathbb{R}_{>0}$.
  2. (b)
@@
  3. (c)
    The integral of the absolute value of $\alpha$ is finite:
    (6) $\rho(\infty)$ $=\int_{0}^{\infty}|\alpha(s)|\,\mathrm{d}s<\infty.$
@@
A scalar-valued kernel, $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}$, is exponentially bounded if there exist $L,\rho\in\mathbb{R}_{>0}$ such that
(7) $|\alpha(t)|$ $\leq Le^{-\rho t}$
@@
Next, we present a few corollaries from the literature about the steady states of (3b)–(4) and their stability.
@@
(8) $0$ $=f(\bar{x},\bar{z}),$ $\bar{z}$ $=\beta(\infty)\odot\bar{r},$ $r(t)$ $=h(\bar{x})m$
where the $i$’th element of $\beta:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{n_{z}}$ is given by (5) in Definition 3.1.
Proof 3.5.
@@
(9) $z(t)$ $=\int_{-\infty}^{t}\alpha(t-s)\odot\bar{r}\,\mathrm{d}s=\int_{-\infty}^{t}\alpha(t-s)\,\mathrm{d}s\odot\bar{r}=\beta(\infty)\odot\bar{r},$
for all $t$.
Corollary 3.6.
The linearization of the system (3b)–(4) around a steady state, $\bar{x}$, satisfying (8) describes the evolution of the deviation variable $X:\mathbb{R}\rightarrow\mathbb{R}^{n_{x}}$:
The matrices $F\in\mathbb{R}^{n_{x}\times n_{x}}$, $G\in\mathbb{R}^{n_{x}\times n_{z}}$, and $H\in\mathbb{R}^{n_{z}\times n_{x}}$ are the Jacobians of the right-hand side function and the delay function evaluated in the steady state:
@@
where $P:\mathbb{C}\rightarrow\mathbb{C}$ is the characteristic function and $I\in\mathbb{R}^{n_{x}\times n_{x}}$ is an identity matrix.
Proof 3.8.
Theorem 4.7 in the paper by Diekmann and Gyllenberg [11] provides conditions under which this result holds, and the assumptions on the system (3b)–(4) ensure that these conditions are satisfied (see Section SM1 in the supplementary material for further details).
4 Main theoretical results
In this section, we present and prove the two main theoretical results of the paper: 1) We can approximate regular kernels arbitrarily well using Erlang mixture approximations (Proposition 4.2), and 2) we can approximate the DDEs (3b)–(4) by ODEs based on Erlang mixture approximations because of convergence results for the solution and steady state of the ODEs and for the roots of the characteristic polynomial as well as a result on the stability of the error of the solution (Theorem 4.4). For simplicity of the presentation, we only consider scalar kernels in Proposition 4.2 on the convergence of Erlang mixture approximations.
First, we define Erlang kernels which constitute the basis functions in the Erlang mixture kernel approximation that we describe afterwards. Figure 1 shows examples of Erlang kernels and Erlang mixture kernels for different rate parameters and coefficients.
@@ 4 Main theoretical results
Proposition 4.2 (Erlang mixture approximation).
Let $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}$ be a scalar-valued regular kernel, and let $\hat{\alpha}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}$ be an Erlang mixture kernel of order $M\in\mathbb{N}_{\geq 0}$ and rate parameter $a\in\mathbb{R}_{>0}$ defined by
(15) $\hat{\alpha}(t)$ $=\sum_{m=0}^{M}c_{m}\ell_{m}(t),$ $c_{m}$ $=\int_{s_{m}}^{s_{m+1}}\alpha(s)\,\mathrm{d}s,$ $s_{m}$ $=m\Delta s,$ $\Delta s$ $=1/a,$
where $c_{m}\in\mathbb{R}$ is the $m$’th coefficient given by the integral of $\alpha$ over a set of intervals with boundaries $s_{m}\in\mathbb{R}_{\geq 0}$ for $m=0,\ldots,M+1$ where the width $\Delta s\in\mathbb{R}_{>0}$ is equal to the inverse of the rate parameter, $a$. Then,
  1. (a)
    $\hat{\alpha}$ is regular,
and the following asymptotic properties hold as $a\rightarrow\infty$ provided that $M/a\rightarrow\infty$ as well.
  1. (b)
    $\hat{\alpha}$ converges uniformly to $\alpha$ for all $t\in\mathbb{R}_{\geq 0}$, i.e.,
    (16) $\hat{\alpha}(t)$ $\rightarrow\alpha(t).$
  2. (c)
    $\hat{\alpha}$ converges weakly to $\alpha$ for all $t\in\mathbb{R}_{\geq 0}$, i.e.,
    (17) $\int_{0}^{t}\hat{\alpha}(s)\,\mathrm{d}s$ $\rightarrow\int_{0}^{t}\alpha(s)\,\mathrm{d}s.$
  3. (d)
    The integral of the absolute difference between $\hat{\alpha}$ and $\alpha$ converges to zero:
    (18) $\int_{0}^{\infty}|\hat{\alpha}(t)-\alpha(t)|\,\mathrm{d}t$ $\rightarrow 0.$
Corollary 4.3.
The Erlang mixture kernel, $\hat{\alpha}$, defined in Proposition 4.2 is exponentially bounded if $\alpha$ is exponentially bounded. Specifically, if $|\alpha(t)|\leq Le^{-\rho t}$ for all $t\in\mathbb{R}_{\geq 0}$ where $L,\rho\in\mathbb{R}_{>0}$, then $|\hat{\alpha}(t)|\leq Le^{-\bar{\rho}t}$ where $\bar{\rho}=\rho(1-e^{-\omega})/\omega\in\mathbb{R}_{>0}$. Furthermore, $0\leq\bar{\rho}<\rho$ and $\bar{\rho}\rightarrow\rho$ as $\omega\rightarrow 0$.